You have 3 projects to consider.

Project A with current and projected future rate of return

Discount Factor at            17%.

Yr

Inflows

Outflows

Netflow

Discounted netflow

Discounted rate

0

-$450,000

-$450,000

-$450,000

1.0000

1

$150,000

$150,000

$128,205

0.8547

2

$150,000

$150,000

$109,577

0.7305

3

$150,000

$150,000

$93,656

0.6244

4

$150,000

$150,000

$80,048

0.5337

5

$150,000

$150,000

$68,417

0.4561

Project B with current and projected future rate of return

Discount Factor   17%

Yr

Inflows

Outflows

Netflow

Discounted netflow

Discounted rate

0

-$400,000

-$400,000

-$400,000

1.0000

1

$0

$0

$0

0.8547

2

$50,000

$50,000

$36,526

0.7305

3

$200,000

$200,000

$124,874

0.6244

4

$300,000

$300,000

$160,095

0.5337

5

$200,000

$200,000

$91,222

0.4561

Project C: with current and projected future rate of return

Discount Factor               17%

Yr

Inflows

Outflows

Netflow

Discounted netflow

Discounted rate

0

-$105,250

-$105,250

-$105,250

1.0000

1

$25,000

$25,000

$21,368

0.8547

2

$25,000

$25,000

$18,263

0.7305

3

$25,000

$25,000

$15,609

0.6244

4

$25,000

$25,000

$13,341

0.5337

answer the following questions.

1.

Project A NPV. You do not include the initial $XXXXX cost as one of the values, because the payment occurs at the beginning of the first period is $29,902

2.

the highest project is A.

3.

Project C NPV. You include the initial $XXXXX cost as one of the values, because the payment occurs at the end of the first period is not $18,180

4.

Project B NPV. You do not include the initial $XXXXX cost as one of the values, because the payment occurs at the beginning of the first period is $13,717

5.

Project B NPV. You include the initial $XXXXX cost as one of the values, because the payment occurs at the end of the first period is $10,869

6.

Project C NPV. You do not include the initial $XXXXX cost as one of the values, because the payment occurs at the beginning of the first period is -$21,271

7.

Project A NPV. You include the initial $XXXXX cost as one of the values, because the payment occurs at the end of the first period is $26,557
Answer #1


Project A NPV=-450000+150000/17%*(1-1/1.17^5)=29901.92

Project B NPV=-400000+0/1.17+50000/1.17^2+200000/1.17^3+300000/1.17^4+200000/1.17^5=12717.03378

Project C NPV=-105250+25000/17%*(1-1/1.17^4)=-36669.1247

the highest project is A.
You have to include the initial cost for all projects